OpenStudy (anonymous):

can i get some help on this guys? please? Solve the equation by graphing x2 - 5x - 24 = 0 A. {4, 6} B. {-2, 12} C. {-4, 6} D. {-3, 8}

OpenStudy (anonymous):

The equation x2 - 5x - 24 = 0 can be factored as (x-8)(x+3) = 0, so we know the solutions are x=8 and x=-3. Here's the solution using the quadratic formula. For ax2+bx+c = 0, the solutions are given by x = (-b ±√(b2-4ac))/(2a) Your equation has a=1, b=-5, c=-24. x = (-(-5) ±√((-5)2 - (4)(1)(-24))/(2*1) = (5 ±√(25 + 96))/2 = (5±√121)/2 = (5±11)/2 x = 16/2 or -6/2 x = 8 or -3 So D

OpenStudy (anonymous):

thank u sooo muchh!!! :)